A popular abdominal exercise is called the "Scissor Abs". Logan has completed 35 full lifts with both legs. He then continues to lie on the floor with his left leg pointing straight up (bottom of foot facing up so that his left leg is perpendicular to the floor) and his right leg resting on the floor. If Logan allows his leg to fall freely until it hits the floor what is the tangential speed of his left foot just before it hits the floor. Assume his leg can be treated as a uniform rod of length 0.95 m long that pivots freely about the hip.
Answer #1

Here, we can use conservation of energy

mgh = 1/2Iw2

where h = 0.95 / 2 = 0.475 m

I = mL2 / 12 + m(L / 2)2 = mL2 / 3 ( parallel axis theorem)

w = v/L

so,

mgh = 1/2 * 1/3 * m * L2 * v2 / L2

cancel all the like terms , we have

v2 = 6gh

v = sqrt (6gh)

v = sqrt ( 6 * 9.8 * 0.475)

v = 5.2848 m/s